OpenStudy (anonymous):

Solve: 5x + 1 = 2x + 7

OpenStudy (anonymous):

Do you have answer choices

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

The answer x=2

OpenStudy (anonymous):

subtract 2x from both sides and then 1 you will end up with x=2

OpenStudy (anonymous):

5(2) + 1= 2(2) + 7 10 + 1= 4 + 7 11 = 11 x= 11

OpenStudy (anonymous):

6. Mrs. Robinson takes a vacation to Maine to pick blueberries. For each bucket of blueberries she picks she gets $8.00. Mrs. Robinson wants to pick enough blueberries to cover her lobster dinner that night which will cost her around $60.00. How many buckets of blueberries will she have to pick to pay for her dinner? *